OpenStudy (anonymous):

Which of these numbers is a solution to the inequality x minus 9 is greater than or equal to nineteen? 9 10 20 28

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

might help to write that out in a more mathical notation

OpenStudy (amistre64):

just test them out ... is 9-9 >= 19? is 10-9 >= 19? is 20-9 >= 19? is 28-9 >= 19?

OpenStudy (anonymous):

28.. because its the only answer that when 9 is subtracted, is greater than or equal to 19

OpenStudy (anonymous):

can you help me to my problem

OpenStudy (anonymous):

l=a+(n-1)d solve for n